# Knowledge Distillation Driven Semantic NOMA for Image Transmission with Diffusion Model

### v1

@@ III-A Teacher Model Training
III-B Student Model Training
@@
IV Diffusion Model-Based Image Refinement
@@ IV-A Preliminary
IV-B Diffusion-Based Image Refinement
In our proposed KDD-SemNOMA scheme, we first leverage the proposed KD-SemNOMA model (as depicted in Section III) to reconstruct images $\mathbf{x}_{\text{init}}=\hat{\mathbf{x}}$ from multi-user non-orthogonal transmissions (Eq. 4). We denote the real high-quality images as $\mathbf{x}$ and the first-stage reconstructed images from the KD-SemNOMA model as $\mathbf{x}_{\text{init}}$. Due to compression losses, channel noise and multi-user interference inherent in the semantic NOMA scenario, $\mathbf{x}_{\text{init}}$ deviates from $\mathbf{x}$, introducing reconstruction errors. To refine $\mathbf{x}_{\text{init}}$ into a high-fidelity approximation of $\mathbf{x}$, we propose designing a posterior distribution $p(\mathbf{x}\mid\mathbf{x}_{\text{init}})$, inspired by the diffusion-based image enhancement strategy in DifFace [32].
@@ IV-B Diffusion-Based Image Refinement
5:Reshape $\mathbf{x}_{\text{init}}\in\mathbb{R}^{B\times N\times C\times H\times W}$ to $\mathbf{x}_{\text{init}}^{\text{flat}}\in\mathbb{R}^{(B\cdot N)\times C\times H\times W}$
@@
8:  Predict noise $\epsilon_{\theta}(x_{t},t)$
9:  Reverse sampling
@@
11:end for
12:Reshape ${\mathbf{x}_{0}}^{\text{flat}}=\{\mathbf{x}_{0}^{(i)}\}_{i=1}^{B\times N}$ to ${\mathbf{x}_{0}}\in\mathbb{R}^{B\times N\times C\times H\times W}$
13:return ${\mathbf{x}_{0}}=\{\mathbf{x}_{0}^{(1)},\dots,\mathbf{x}_{0}^{(N)}\}$
V Training and Evaluation
V-A Simulation Setup
V-A1 Datasets
@@
V-A2 Training and Inference Details
We train the KD-SemNOMA framework following Algorithm 1. Specifically, we employ the AdamW optimizer [38] with a learning rate of $1\times 10^{-4}$ and a weight decay of $1\times 10^{-4}$. Both the teacher and student models are trained using the batch size of 32. To mitigate overfitting, an early stopping mechanism is implemented, terminating training when no improvement in validation performance is observed over 10 consecutive epochs. During student model training, the loss weights are set to $\lambda_{1}=10$, $\lambda_{2}=100$, and $\lambda_{3}=1$ to balance the contributions of the restoration loss and distillation terms, as defined in Section III. During inference, KD-SemNOMA relies solely on the student model, incurring no additional computational overhead compared to the proposed SemNOMA model as detailed in Section II-C. For DDIM inference, the first-stage output $\mathbf{x}_{\text{init}}=\hat{\mathbf{x}}$ is refined through iterative denoising (Eq. 23), using $\eta=0.5$ and $T^{\prime}=200$ steps to maximize image perceptual performance.
For datasets of different resolutions, the encoder and decoder network parameters of the proposed ConvNeXt-based DeepJSCC are detailed in Table I for CIFAR-10 and Table II for FFHQ-256. The proposed framework and all comparative methods are implemented in PyTorch and executed on an NVIDIA 4090 GPU server.
TABLE I: Parameters setting of DeepJSCC encoder and decoder for the image resolution of 32×32
@@ V-A2 Training and Inference Details
TABLE II: Parameters setting of DeepJSCC encoder and decoder for the image resolution of 256×256
@@
V-A3 Benchmark Schemes
  • JPEG+LDPC+QAM+SIC: In this case, the conventional scheme performs the source coding and channel coding separately. Joint photographic experts group (JPEG) and low-density parity-check (LDPC) are employed for the source and channel coding. The LDPC code employs $3/4$ code rate. The modulation scheme is 4 quadrature amplitude modulation (QAM). For multi user detection at the receiver, the successive interference cancellation (SIC) algorithm [39] is employed.
  • SemNOMA-CGAN: In this case, borrow from GAN decomposition method [31], we combine the proposed SemNOMA with conditional GAN (CGAN) network to enhance the perceptual quality of the multi-user restored images. We utilize the U-Net [40] based generator and use the combination of MAE loss, learned perceptual image patch similarity (LPIPS) loss, and GAN loss to train the CGAN network.
V-A4 Performance Metrics
@@
V-B Performance of Knowledge Distillation-Based SemNOMA for Image Transmission
@@
We evaluate the proposed KD-SemNOMA framework for image transmission, comparing it against baselines (i.e., JPEG+LDPC+QAM+SIC, DeepJSCC-NOMA) and semantic orthogonal transmission schemes (i.e., SemOMA) on CIFAR-10 and FFHQ-256 datasets under AWGN and Rayleigh fading channels, with 2 UE and compression ratios $\rho$ of 1/6 and 1/3 for CIFAR-10, and 1/48 for FFHQ-256. The ConvNeXt-based KD-SemNOMA, optimized via Algorithm 1, leverages knowledge distillation stragety to enhance pixel-level fidelity (i.e., PSNR and SSIM), as detailed below.
We first conduct simulation comparison between the proposed KD-SemNOMA and the baseline scheme. As shown in Fig. 6, we evaluate the image reconstruction performance in terms of PSNR and SSIM on the CIFAR-10 dataset under the AWGN channel with a compression ratio $\rho$ of $1/6$ and $1/3$ across various $\gamma$ conditions. The results indicate that the JPEG+LDPC+QAM+SIC method suffers from a cliff effect at low SNR, leading to a significant degradation in image reconstruction. In our simulations, we employs non-equal power allocation (power factors of 0.7 and 0.3 for the two users) in JPEG+LDPC+QAM+SIC to enable feasible SIC decoding, yet it still fails to overcome the interference-limited performance floor. Additionally, the conventional NOMA scheme with SIC detection algorithm fails to effectively separate signals and mitigate interference in equal power allocation scenarios, resulting in poor PSNR and SSIM performance. In contrast, the proposed deep learning-based multi-user image transmission framework performs robustly at low SNR and demonstrates strong adaptability to arbitrary power allocation scenarios. Compared to the baseline DeepJSCC-NOMA framework based on the ResNet architecture, our proposed SemNOMA scheme, leveraging the ConvNeXt network, outperforms the baseline across various $\gamma$ conditions, validating the effectiveness of the ConvNeXt network in extracting deep image features.
@@
(a) LPIPS vs. SNR (AWGN)
(b) FID vs. SNR (AWGN)
(c) LPIPS vs. SNR (Rayleigh)
(d) FID vs. SNR (Rayleigh)
We also conduct simulations on the high-resolution FFHQ-256 dataset to verify the compatibility of the proposed scheme with different resolution scenarios. As shown in Fig. 8, we evaluate the image reconstruction PSNR and SSIM performance of different schemes under AWGN and Rayleigh channel with a compression ratio of $1/48$ across various SNR conditions. Fig. 8(a) and Fig. 8(b) shows the simulation in AWGN channel. Compared to the baseline DeepJSCC-NOMA scheme, the proposed ConvNeXt-based framework SemNOMA demonstrates a more pronounced advantage in high-resolution scenarios, achieving a PSNR gain of $0.3\sim 0.4$dB, with further improvements of $0.3\sim 0.5$dB after knowledge distillation optimization in KD-SemNOMA. The SSIM metric also shows consistent improvement. Compared to the orthogonal transmission scheme SemOMA, the proposed non-orthogonal scheme KD-SemNOMA achieves a performance gain of approximately $0.3\sim 0.5$dB at the same overhead. Fig. 8(c) and Fig. 8(d) shows the simulation in Rayleigh channel. Compared to the baseline DeepJSCC-NOMA, the proposed SemNOMA framework achieves a performance gain of $1\sim 1.5$dB, with knowledge distillation optimization providing an additional gain of approximately 0.40.4dB in KD-SemNOMA. In terms of PSNR, the proposed KD-SemNOMA performs slightly worse than the orthogonal scheme SemOMA with the same transmission overhead in the [0,6][0,6]dB range but outperforms it in the [6,20][6,20]dB range. For the SSIM metric, the proposed scheme consistently outperforms SemOMA. Overall, in high resolution and high compression ratio cases, our scheme demonstrates significant advantages over both the non-orthogonal baseline and orthogonal schemes, confirming the effectiveness of the proposed framework.
TABLE III: Effectiveness of applying CrossKD at different positions in DeepJSCC decoder at $\gamma=10$ dB
@@
As described in Section III, CrossKD transfers the intermediate features of the $j$-th layer from the student model decoder to the corresponding position of the teacher model decoder for forward propagation, enhancing feature alignment in knowledge distillation (Algorithm 1). To verify the effect of the choice of $j$ on the distillation performance, we conduct the following experiments. Cases include A32C3 (AWGN, CIFAR-10, $\rho=1/3$), R32C3 (Rayleigh, CIFAR-10, $\rho=1/3$), A256C48 (AWGN, FFHQ-256, $\rho=1/48$), and R256C48 (Rayleigh, FFHQ-256, $\rho=1/48$). As shown in Table III, the first row represents the student model trained without teacher model distillation, using only MAE loss. For A32C3, $j=4$ (upsampling layer) yields the optimal PSNR/SSIM (33.18/0.973), as upsampling features align well with AWGN’s stable conditions. For R32C3, $j=2$ (AF-Module layer) is the best (29.58/0.927), leveraging AF-Module’s adaptation to Rayleigh fading’s variability. In FFHQ-256 cases, $j=5$ performs best (29.24/0.856) for A256C48 while $j=4$ performs best (27.52/0.818) for R256C48. These results highlight CrossKD’s ability to optimize layer-specific feature transfer, enhancing decoding performance across diverse datasets and channel conditions.
V-C Performance of Diffusion-Based Image Refinement for SemNOMA
@@
Simulations are conducted under AWGN and Rayleigh fading channels with the compression ratio $\rho=1/48$, across SNR $\gamma$ ranges from -5 to 20 dB. We compare the proposed KDD-SemNOMA scheme against CGAN based image refinement SemNOMA-CGAN, and the one-stage schemes (KD-SemNOMA, SemNOMA, DeepJSCC-NOMA). As shown in Fig. 9, KDD-SemNOMA achieves superior LPIPS and FID across all SNR levels, with particularly notable improvements in the low SNR range of [-5, 5] dB. For instance, at $\gamma$ = 0dB under AWGN channel, KDD-SemNOMA significantly reduces the LPIPS and FID values compared with the one-stage method, indicating that the utilize of the diffusion model greatly improves the perceptual quality of the received image. At the same time, compared with SemNOMA-CGAN, KDD-SemNOMA reduces the LPIPS and FID values by approximately $23\%$ and $36\%$, respectively, reflecting the image generation advantages of the diffusion model over the CGAN model and its robustness to severe channel noise.
@@
Original DeepJSCC-NOMA SemNOMA KD-SemNOMA SemNOMA-CGAN KDD-SemNOMA
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Original DeepJSCC-NOMA SemNOMA KD-SemNOMA SemNOMA-CGAN KDD-SemNOMA
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
We provide examples of reconstructed images from FFHQ-256 datasets in Fig. 10 and Fig. 11, showcasing visual quality at $\gamma=0$ dB under AWGN and Rayleigh fading channels, respectively. While KD-SemNOMA and SemNOMA-CGAN achieve competitive PSNR and SSIM for initial reconstructions $\mathbf{x}_{\text{init}}=\hat{\mathbf{x}}_{i}$, their images suffer from noise-induced degradation at $\gamma=0$ dB, losing semantic details such as facial contours. SemNOMA-CGAN, relying on GAN-based generation, often introduces artifacts and blurred textures, particularly under Rayleigh fading, due to mode collapse in low-SNR conditions. In contrast, KDD-SemNOMA leverages error contraction in the forward diffusion process (Eq. 24) and iterative denoising with diffusion priors (Eq. 25), achieving superior LPIPS and FID scores. Visualizations demonstrate that KDD-SemNOMA restores finer facial details, such as hair textures and eye clarity, outperforming SemNOMA-CGAN by producing more natural and perceptually coherent images [32].
V-D Ablation Study
To validate the effectiveness of the knowledge distillation optimization scheme in enhancing model performance, we conduct ablation studies using following datasets (CIFAR-10, FFHQ-256), compression ratios (1/3, 1/48), and channel conditions (AWGN, Rayleigh) at $\gamma=10$ dB. Results are presented in Table IV, with cases defined as follows: A32C3 (AWGN, CIFAR-10, $\rho=1/3$), R32C3 (Rayleigh, CIFAR-10, $\rho=1/3$), A256C48 (AWGN, FFHQ-256, $\rho=1/48$), and R256C48 (Rayleigh, FFHQ-256, $\rho=1/48$). The first row represents the baseline where only reconstruction loss is used, indicating that the student model is trained without distillation from the teacher model. The subsequent rows demonstrate the performance of the student model trained with different distillation algorithms. Specifically, the second and third rows confirm that both FA and CrossKD distillation improve model performance. The final row indicates that the best performance is achieved when FA and CrossKD are applied simultaneously. Notably, performance gains are consistent across datasets and channels. These findings align with results in Fig. 6 and Fig. 7 for CIFAR-10 and Fig. 8 for FFHQ-256.
TABLE IV: Effects of different distillation components on the performance of the student model at $\gamma=10$ dB

### v2

@@ IV-A Teacher Model Training
IV-B Student Model Training
@@
IV-C Discussion
In summary, our KD-SemNOMA framework establishes a comprehensive teacher-student interaction protocol that enables effective knowledge transfer across multiple levels that effectively bridges the performance gap between orthogonal and non-orthogonal transmission systems. The key aspects of our approach include:
  1. 1.
    Multi-level Knowledge Transfer: During training, we implement knowledge distillation at both feature-level (through Feature Affinity loss) and prediction-level (through CrossKD loss), ensuring comprehensive guidance from the teacher model to the student model.
  2. 2.
    Efficient Training Protocol: The progressive training strategy with teacher model initialization provides a strong foundation, while the adaptive loss balancing mechanism ensures stable optimization throughout the training process.
  3. 3.
    Computational Efficiency: During inference, only the lightweight student model is deployed with no teacher-student interaction, ensuring no additional computational overhead compared to baseline SemNOMA system.
The performance improvement achieved by our method is particularly significant in challenging low-SNR scenarios, where the knowledge distillation effectively mitigates the impact of both channel noise and multi-user interference. While perfect knowledge transfer cannot be guaranteed due to the fundamental constraints of non-orthogonal transmission, our experimental results in Section VI-B demonstrate that the student model consistently outperforms baseline methods without knowledge distillation.
V Diffusion Model-Based Image Refinement
@@ V-A Preliminary
V-B Diffusion-Based Image Refinement
In our proposed KDD-SemNOMA scheme, we first leverage the proposed KD-SemNOMA model (as depicted in Section IV) to reconstruct images $\mathbf{x}_{\text{init}}=\hat{\mathbf{x}}$ from multi-user non-orthogonal transmissions (Eq. 4). We denote the real high-quality images as $\mathbf{x}$ and the first-stage reconstructed images from the KD-SemNOMA model as $\mathbf{x}_{\text{init}}$. Due to compression losses, channel noise and multi-user interference inherent in the semantic NOMA scenario, $\mathbf{x}_{\text{init}}$ deviates from $\mathbf{x}$, introducing reconstruction errors. To refine $\mathbf{x}_{\text{init}}$ into a high-fidelity approximation of $\mathbf{x}$, we propose designing a posterior distribution $p(\mathbf{x}\mid\mathbf{x}_{\text{init}})$, inspired by the diffusion-based image enhancement strategy in DifFace [43].
@@ V-B Diffusion-Based Image Refinement
5:Reshape $\mathbf{x}_{\text{init}}=\{\mathbf{x}_{\text{init}}^{(i)}\}_{i=1}^{N}$ to $\mathbf{x}_{\text{init}}^{\text{flat}}\in\mathbb{R}^{(B\cdot N)\times C\times H\times W}$
@@
8:  Reverse sampling
@@
10:end for
11:Reshape ${\mathbf{x}_{0}}^{\text{flat}}\in\mathbb{R}^{(B\cdot N)\times C\times H\times W}$ to ${\mathbf{x}_{0}}=\{\mathbf{x}_{0}^{(i)}\}_{i=1}^{N}$
12:return ${\mathbf{x}_{0}}=\{\mathbf{x}_{0}^{(1)},\dots,\mathbf{x}_{0}^{(N)}\}$
VI Training and Evaluation
VI-A Simulation Setup
VI-A1 Datasets
@@
VI-A2 Benchmark Schemes
  • BPG+LDPC+QAM+SIC: In this case, the conventional scheme performs the source coding and channel coding separately. Better Portable Graphics (BPG) and low-density parity-check (LDPC) are employed for the source and channel coding. The $1/2$ code rate LDPC code is employed with information block length of 4096 bits and codeword length of 8192 bits. The modulation scheme is 4 quadrature amplitude modulation (QAM). For multi-user detection at the receiver, the successive interference cancellation (SIC) algorithm [49] is employed.
  • SemNOMA-CGAN: In this case, borrow from GAN decomposition method [42], we combine the proposed SemNOMA with conditional GAN (CGAN) network to enhance the perceptual quality of the multi-user restored images. We utilize the U-Net [50] based generator and use the combination of MAE loss, learned perceptual image patch similarity (LPIPS) loss, and GAN loss to train the CGAN network. Detailed training implementation of SemNOMA-CGAN is provided in Section VI-A3.
  • SemNOMA-DiffBIR: Build upon the two-stage framework of DiffBIR [51], we integrate the SemNOMA with the pretrained diffusion-based generation module from DiffBIR, which introduces the IRControlNet that leverages Stable Diffusion [52] prior for realistic image restoration. For inference, LLaVA-7B [53] is applied to the SemNOMA initial output to generate a descriptive caption, which is then concatenated with a fixed high-quality positive prompt. The resulting text prompt, together with the initial output image as condition, is fed into the Stable Diffusion model equipped with IRControlNet for generative sampling.
VI-A3 Training and Inference Details
We train the KD-SemNOMA framework following Algorithm 1. Specifically, we employ the AdamW optimizer [54] with a learning rate of $1\times 10^{-4}$ and a weight decay of $1\times 10^{-4}$. Both the teacher and student models are trained using the batch size of 32. To mitigate overfitting, an early stopping mechanism is implemented, terminating training when no improvement in validation performance is observed over 10 consecutive epochs. During student model training, the loss weights are set to $\lambda_{1}=10$, $\lambda_{2}=100$, and $\lambda_{3}=1$ to balance the contributions of the restoration loss and distillation terms, as defined in Section IV. During inference, KD-SemNOMA relies solely on the student model, incurring no additional computational overhead compared to the proposed SemNOMA model as detailed in Section III-C. For DDIM inference, the first-stage output $\mathbf{x}_{\text{init}}=\hat{\mathbf{x}}$ is refined through iterative denoising (Eq. 23), using $\eta=0.5$ and $T^{\prime}=200$ steps to maximize image perceptual performance. For the SemNOMA-CGAN baseline, followed [42], we employ a CGAN with a U-Net based generator and a PatchGAN discriminator. The CGAN is trained using the outputs from the pre-trained KD-SemNOMA model as conditional input. During training, the KD-SemNOMA weights remain frozen while the CGAN components are optimized through AdamW optimizer over 100 epochs with a composite loss function where the weights are set to 1 for MAE, 0.1 for LPIPS, and $1\times 10^{-4}$ for the adversarial loss.
For datasets of different resolutions, the encoder and decoder network parameters of the proposed ConvNeXt-based DeepJSCC are detailed in Table II for CIFAR-10 and Table III for FFHQ-256. The proposed framework and all comparative methods are implemented in PyTorch and executed on an NVIDIA 4090 GPU server.
TABLE II: Parameters setting of DeepJSCC encoder and decoder for the image resolution of 32×32
@@ VI-A3 Training and Inference Details
TABLE III: Parameters setting of DeepJSCC encoder and decoder for the image resolution of 256×256
@@
VI-A4 Performance Metrics
@@
VI-B Performance of Knowledge Distillation-Based SemNOMA for Image Transmission
@@
We evaluate the proposed KD-SemNOMA framework for image transmission, comparing it against baselines (i.e., BPG+LDPC+QAM+SIC, DeepJSCC-NOMA) and semantic orthogonal transmission schemes (i.e., SemOMA) on CIFAR-10 and FFHQ-256 datasets under AWGN and Rayleigh fading channels, with 2 UE and compression ratios $\rho$ of 1/6 and 1/3 for CIFAR-10, and 1/48 for FFHQ-256. The ConvNeXt-based KD-SemNOMA, optimized via Algorithm 1, leverages knowledge distillation stragety to enhance pixel-level fidelity (i.e., PSNR and SSIM), as detailed below.
We first conduct simulation comparison between the proposed KD-SemNOMA and the baseline scheme. As shown in Fig. 6, we evaluate the image reconstruction performance in terms of PSNR and SSIM on the CIFAR-10 dataset under the AWGN channel with a compression ratio $\rho$ of $1/6$ and $1/3$ across various $\gamma$ conditions. The results indicate that the BPG+LDPC+QAM+SIC method suffers from a cliff effect at low SNR, leading to a significant degradation in image reconstruction. In our simulations, we employs non-equal power allocation (power factors of 0.8 and 0.2 for the two users) in BPG+LDPC+QAM+SIC to enable feasible SIC decoding, yet it still fails to overcome the interference-limited performance floor. Additionally, the conventional NOMA scheme with SIC detection algorithm fails to effectively separate signals and mitigate interference in equal power allocation scenarios, resulting in poor PSNR and SSIM performance. In contrast, the proposed deep learning-based multi-user image transmission framework performs robustly at low SNR and demonstrates strong adaptability to arbitrary power allocation scenarios. Compared to the baseline DeepJSCC-NOMA framework based on the ResNet architecture, our proposed SemNOMA scheme, leveraging the ConvNeXt network, outperforms the baseline across various $\gamma$ conditions, validating the effectiveness of the ConvNeXt network in extracting deep image features.
@@
(a) CrossKD Loss vs. Epoch
(b) FA Loss vs. Epoch
(c) MAE Loss vs. Epoch
(d) Total Loss vs. Epoch
Figure 9: Visualization for the variation of statistics during training under cases of A256C48 and R256C48.
We also conduct simulations on the high-resolution FFHQ-256 dataset to verify the compatibility of the proposed scheme with different resolution scenarios. As shown in Fig. 8, we evaluate the image reconstruction PSNR and SSIM performance of different schemes under AWGN and Rayleigh channel with a compression ratio of $1/48$ across various SNR conditions. Fig. 8(a) and Fig. 8(b) shows the simulation in AWGN channel. Compared to the baseline DeepJSCC-NOMA scheme, the proposed ConvNeXt-based framework SemNOMA demonstrates a more pronounced advantage in high-resolution scenarios, achieving a PSNR gain of $0.3\sim 0.4$dB, with further improvements of $0.3\sim 0.5$dB after knowledge distillation optimization in KD-SemNOMA. The SSIM metric also shows consistent improvement. Compared to the orthogonal transmission scheme SemOMA, the proposed non-orthogonal scheme KD-SemNOMA achieves a performance gain of approximately $0.3\sim 0.5$dB at the same overhead. Notably, the traditional BPG+LDPC+QAM+SIC scheme exhibits the typical cliff effect of separation-based approaches, where the reconstruction quality drops sharply when the channel quality falls below a certain threshold (around 5-6 dB SNR). In contrast, all semantic communication-based schemes demonstrate graceful performance degradation in low-SNR regimes. Fig. 8(c) and Fig. 8(d) shows the simulation in Rayleigh channel. Compared to the baseline DeepJSCC-NOMA, the proposed SemNOMA framework achieves a performance gain of $1\sim 1.5$dB, with knowledge distillation optimization providing an additional gain of approximately 0.40.4dB in KD-SemNOMA. In terms of PSNR, the proposed KD-SemNOMA performs slightly worse than the orthogonal scheme SemOMA with the same transmission overhead in the [0,6][0,6]dB range but outperforms it in the [6,20][6,20]dB range. For the SSIM metric, the proposed scheme consistently outperforms SemOMA. The BPG+LDPC+QAM+SIC baseline shows even more severe performance degradation under Rayleigh fading conditions: the cliff effect occurs at higher SNR levels (around 10-11 dB) and decoding fails completely below these thresholds. Overall, in high resolution and high compression ratio cases, our scheme demonstrates significant advantages over both the non-orthogonal baseline and orthogonal schemes, confirming the effectiveness of the proposed framework.
TABLE IV: Effectiveness of applying CrossKD at different positions in DeepJSCC decoder at $\gamma=10$ dB
@@
As described in Section IV, CrossKD transfers the intermediate features of the $j$-th layer from the student model decoder to the corresponding position of the teacher model decoder for forward propagation, enhancing feature alignment in knowledge distillation (Algorithm 1). To verify the effect of the choice of $j$ on the distillation performance, we conduct the following experiments. Cases include A32C3 (AWGN, CIFAR-10, $\rho=1/3$), R32C3 (Rayleigh, CIFAR-10, $\rho=1/3$), A256C48 (AWGN, FFHQ-256, $\rho=1/48$), and R256C48 (Rayleigh, FFHQ-256, $\rho=1/48$). As shown in Table IV, the first row represents the student model trained without teacher model distillation, using only MAE loss. For A32C3, $j=4$ (upsampling layer) yields the optimal PSNR/SSIM (33.18/0.973), as upsampling features align well with AWGN’s stable conditions. For R32C3, $j=2$ (AF-Module layer) is the best (29.58/0.927), leveraging AF-Module’s adaptation to Rayleigh fading’s variability. In FFHQ-256 cases, $j=5$ performs best (29.24/0.856) for A256C48 while $j=4$ performs best (27.52/0.818) for R256C48. These results highlight CrossKD’s ability to optimize layer-specific feature transfer, enhancing decoding performance across diverse datasets and channel conditions.
To validate the effectiveness of our knowledge distillation based training strategy, we conducted a comprehensive analysis of the loss convergence behavior during training. As shown in Fig. 9, both distillation losses ($\mathcal{L}_{\mathrm{FA}}$ and $\mathcal{L}_{\text{CrossKD}}$) exhibit smooth and stable convergence patterns across different channel conditions, indicating effective knowledge transfer from the teacher model to the student model. Importantly, the distillation losses and the reconstruction loss ($\mathcal{L}_{\text{MAE}}$) demonstrate complementary optimization behavior without conflicting trends, suggesting that the knowledge distillation process enhances rather than interferes with the primary reconstruction task. The overall loss $\mathcal{L}_{S}$ shows consistent and stable convergence throughout the training process, further validating the robustness of our multi-loss optimization framework and the effectiveness of the knowledge distillation approach in improving the student model’s performance in non-orthogonal transmission scenarios.
VI-C Performance of Diffusion-Based Image Refinement for SemNOMA
(a) LPIPS vs. SNR (AWGN)
(b) FID vs. SNR (AWGN)
(c) LPIPS vs. SNR (Rayleigh)
(d) FID vs. SNR (Rayleigh)
@@ VI-C Performance of Diffusion-Based Image Refinement for SemNOMA
Simulations are conducted under AWGN and Rayleigh fading channels with the compression ratio $\rho=1/48$, across SNR $\gamma$ ranges from -5 to 20 dB. We compare the proposed two-stage KDD-SemNOMA scheme against other two-stage refinement methods (including Stable Diffusion based SemNOMA-DiffBIR and CGAN based SemNOMA-CGAN) and one-stage baseline schemes (KD-SemNOMA, SemNOMA, DeepJSCC-NOMA). As shown in Fig. 10, KDD-SemNOMA achieves superior LPIPS and FID across all SNR levels, with particularly notable improvements in the low SNR range of [-5, 5] dB. For instance, at $\gamma$ = 0dB under AWGN channel, KDD-SemNOMA significantly reduces the LPIPS and FID values compared with the one-stage method, indicating that the utilize of the diffusion model greatly improves the perceptual quality of the received image. At the same time, compared with SemNOMA-CGAN, KDD-SemNOMA reduces the LPIPS and FID values by approximately $23\%$ and $36\%$, respectively, reflecting the image generation advantages of the diffusion model over the CGAN model and its robustness to severe channel noise. Compared with SemNOMA-DiffBIR, which also utilizes the powerful generative capability of diffusion models, the two schemes adopt distinct refinement mechanisms: SemNOMA-DiffBIR extracts both textual and visual features from the initial estimate and uses them as conditional guidance to steer the Stable Diffusion denoising process, whereas KDD-SemNOMA first adds noise to the initial estimate to reduce reconstruction error and then performs reverse denoising to generate high-quality images. This enables KDD-SemNOMA to achieve superior LPIPS and FID performance across different SNR regimes.
@@
We provide examples of reconstructed images from FFHQ-256 datasets in Fig. 11 and Fig. 12, showcasing visual quality at $\gamma=0$ dB under AWGN and Rayleigh fading channels, respectively. While the one-stage method KD-SemNOMA achieve competitive PSNR and SSIM for initial reconstructions $\mathbf{x}_{\text{init}}=\hat{\mathbf{x}}_{i}$, its images suffer from noise-induced degradation at $\gamma=0$ dB, losing semantic details such as facial contours. The two-stage refinement schemes generally improve perceptual quality. However, the GAN-based generation method SemNOMA-CGAN often introduces artifacts and blurred textures due to mode collapse in low-SNR conditions, particularly under Rayleigh fading. In comparison, both diffusion-based two-stage methods, SemNOMA-DiffBIR and KDD-SemNOMA, recover finer semantic details and exhibit enhanced visual coherence. Nevertheless, KDD-SemNOMA more faithfully preserves original texture details and facial features, benefiting from its targeted error contraction in the forward diffusion process (Eq. 24) and iterative denoising guided by diffusion priors (Eq. 25).
VI-D Ablation Study
To validate the effectiveness of the knowledge distillation optimization scheme in enhancing model performance, we conduct ablation studies using following datasets (CIFAR-10, FFHQ-256), compression ratios (1/3, 1/48), and channel conditions (AWGN, Rayleigh) at $\gamma=10$ dB. Results are presented in Table V, with cases defined as follows: A32C3 (AWGN, CIFAR-10, $\rho=1/3$), R32C3 (Rayleigh, CIFAR-10, $\rho=1/3$), A256C48 (AWGN, FFHQ-256, $\rho=1/48$), and R256C48 (Rayleigh, FFHQ-256, $\rho=1/48$). The first row represents the baseline where only reconstruction loss is used, indicating that the student model is trained without distillation from the teacher model. The subsequent rows demonstrate the performance of the student model trained with different distillation algorithms. Specifically, the second and third rows confirm that both FA and CrossKD distillation improve model performance. The final row indicates that the best performance is achieved when FA and CrossKD are applied simultaneously. Notably, performance gains are consistent across datasets and channels. These findings align with results in Fig. 6 and Fig. 7 for CIFAR-10 and Fig. 8 for FFHQ-256.
TABLE V: Effects of different distillation components on the performance of the student model at $\gamma=10$ dB